OpenStudy (anonymous):

Wes owns a shop where he sells souvenirs. He ordered a shipment of boxes of huckleberry chocolates. Only threefourths of his order arrived. At that time he had 5 boxes left. Added to the new boxes, he had a total of 128 boxes of the chocolates. a. Solve the equationthree fourths b plus five to find the number of boxes b that were supposed to be in the shipment. Explain how you solved the equation. b. Check the solution to the equation from part a. Explain your method. (2 points)

OpenStudy (anonymous):

total ordered = b (3/4)b + 5 = 128 (3/4)b = 128 - 5 = 123 b = (4 x 123)/3

OpenStudy (anonymous):

Thanks!!!